# Reducing classical communication costs in multiplexed quantum repeaters using hardware-aware quasi-local policies

### v2

@@ VI.2 Figures of merit and repeater chain design for existing memory platforms
Let us now address the following question. Given a particular number $n_{ch}$ of channels for multiplexing, and given other physical parameters corresponding to channel losses, quantum memory coherence times, etc., what is the optimal number of repeater nodes for a linear quantum repeater chain spanning a certain distance, and what would be the rate and fidelity of end-to-end entanglement distribution for this optimal setting? Intuitively, we do expect such an optimal number of nodes to exist, because as the number of nodes increases, even though the individual links become smaller and thus $p_{\ell}$ and $m^{\star}$ both increase, the non-deterministic nature of entanglement swapping will at some point begin to adversely affect the waiting time as the number of nodes increases.
@@
Figure 12: Average waiting time (top), including classical communication overheads, and average fidelity of the end-to-end link (bottom) for entanglement distribution along a 100 km repeater chain with nine channels with rare-Earth ($\text{Pr}^{3+}$ ion) (left) and diamond vacancy (right) based memories using the FN swap-asap policy. Even though the $T_{2}$ coherence times and elementary link probabilities are relatively low, multiplexed policies provide reasonable communication rates and end-to-end link fidelities. Using three to five repeaters between the end nodes seems to be optimal for both memories in terms of waiting times.
@@
To answer this question, let us set the end-to-end distance to be 100 km. We also use the FN swap-asap policy for illustrative purposes, and we also include the classical communication overheads, but we leave out distillation-based policies. We consider both deterministic and non-deterministic entanglement swapping success probabilities ($p_{sw}=1.0,p_{sw}=0.5$) and set the number of channels for multiplexing to be $n_{ch}=9$. The other physical parameters are also chosen to reflect the proposed experimental implementation above. For the quantum memories, we choose two state-of-the-art quantum memory platforms for our analysis, a rare-Earth metal based memory ($\rm{Pr}^{3+}$ ions) and a diamond vacancy based memory.
@@
We plot the average waiting time and fidelity in Fig. 12. We tabulate the values for the fidelity and entanglement distribution rate for the two memory platforms in the optimal setting, along with the hardware parameters, in Table 2. The optimal number of nodes in terms of the waiting time seems to be around five to seven for both memory platforms, which have $T_{2}$ times on the order of a few milliseconds. From the minimum value of the waiting time, we anticipate entanglement distribution rates ranging from a few tens of Hertz to a few hundred Hertz, with an average end-to-end link fidelity of 60-70%. We also see that as the swapping success probability is increased, especially close to the optimal choice of number of repeaters, there is a substantial decrease in the waiting times, although the end-to-end fidelity does not substantially improve. This is also a artifact of choosing the farthest neighbor policy which prioritizes forming long links. Also, the results of both the platforms indicate that the end-to-end link fidelity decreases with increasing number of nodes.
@@
Since, the elementary links have (nearly) unit fidelity for the rare-Earth metal-based memory, distillation is detrimental as shown by the results in Fig. 10. For the diamond vacancy-based memory, we also find that for the network parameters under consideration distillation is in fact detrimental. This could be attributed to the higher CC overheads compared to the small increase in fidelity that distillation offers in this case. For example, the maximum rates available (for optimal number of repeaters) with deterministic entanglement swapping and using FN-based distill-swap policy is around 10 Hz, which is at least ten times smaller than the rates available without distillation.
Memory $T_{2}$ $\eta$ $f_{0}$ $p_{\ell}$ $m^{\star}$ $m_{0}$ $R$ $f$
Rare-Earth 1.2 ms 0.69 0.98 0.046 9 0 160 Hz 0.70
Diamond 13 ms 0.20 0.89 0.004 104 24 8 Hz 0.56
Table 2: Hardware parameters for two quantum memory platforms (rare-earth ion and diamond vacancy), and their respective optimal entanglement distribution rates ($R$) and end-to-end fidelities ($f$), for a 100 km repeater chain, based on the data in Fig. 12. $T_{2}$ is the coherence time of the memory, $\eta$ is the absorption efficiency (Debye–Waller factor), and $f_{0}$ is the fidelity of a fresh elementary link. Entanglement swapping is performed using linear optics hence the success probability is 50%. Also shown are the parameters $p_{\ell}$, $m^{\star}$, and $m_{0}$ of our theoretical model.
VII Conclusions and directions for future work
@@
In this paper, we have taken steps in these directions. We have presented practical, quasi-local multiplexing-based policies for long-distance entanglement distribution using quantum repeaters with multiple memories. We call these policies farthest neighbour (FN) swap-asap and strongest neighbour (SN) swap-asap, adapting swap-as-soon-as-possible (swap-asap) policies to multiplexing-based linear networks. These policies go beyond fully local policies, such that the nodes have knowledge of states of the other nodes in the chain, but not necessarily full, global knowledge.
@@
Data availability
The data generated through the simulations performed in this work and the experimental data that support the findings of this study are available on request from the corresponding author.
Code availability
The custom code developed to perform the simulations in this work that support the findings of this study are available on request from the corresponding author.
Author contributions
SH, SK, HL, and PB conceived the main idea of the work. SK, PB, and SH developed the theory and performed the computations. SK and SH performed the analytical calculations. XC and KC performed the experiments supervised by CW. BK encouraged SH to investigate the role of classical communication. CW, HL, BK and SK supervised the findings of this work. All authors discussed the results and contributed to the writing of the manuscript.
Competing interests
The authors declare no competing interests.
Acknowledgements.
Appendix A Properties of our multiplexing policies
Figure 13: Average waiting time and average age of the youngest end-to-end link for a repeater chain with $n=7$ nodes, according to the multiplexed (SN and FN swap-asap) and non-multiplexed (parallel swap-asap) policies. The multiplexed versions can enormously reduce average waiting time compared to parallel swap-asap, and at the same time increase end-to-end link fidelities. SN is the better choice for fidelity enhancement, while FN reduces the waiting time more significantly.
In Fig. 13, and for all of the plots shown in this paper, the parameter values that we choose are based on what is computationally feasible, but they are also based on what we expect with actual hardware, as we outline in Sec. VI.1. The choice of network parameters such as $p_{\ell},m^{\star}$ etc. in specific examples is also motivated by the intent to show the most prominent and overarching trends.
Figure 14: Average waiting time and average age of the youngest end-to-end link for a repeater chain with $n=7$ nodes, according to the FN, SN, and random swap-asap multiplexing policies. SN and FN both outperform the random policy, indicating the benefit of using knowledge of the network state. SN is the better choice for fidelity enhancement, while FN reduces the waiting time more significantly.
Figure 15: Heat map of average waiting time (left) (denoted $wt\right>$) and ratio of the average age of youngest end-to-end link to the cutoff (right) as a function of elementary link success probability and number of channels. Contours of fixed values run along diagonals from the top-left to the bottom-right, showing that lower link probabilities can be compensated for by increasing the number of channels. Both the average waiting time and the average age of the youngest end-to-end link reduce as the number of channels increases, showing the usefulness of multiplexing. For the waiting time results we use the FN policy, and for the average age results we use the SN policy.
Figure 16: Heat map of average waiting time (left) (denoted by $wt\right>$) and ratio of the average age of youngest end-to-end link to the cutoff (right) as a function of link cutoff age and number of channels. Contours of fixed values run along diagonals from the top-left to the bottom-right, showing that lower cutoffs, which translates to memories with smaller coherence times, can be compensated for by increasing the number of channels. For waiting time results we use the FN policy and for average age results we use SN policy.
Of course, FN, SN, and random are not the only three possible options for multiplexing-based policies. For example, one could develop a ranking scheme based on a combination of age and distance, or one could perform a search over all possible pairings of links, looking for the optimal pairings in terms of some figure of merit (such as the sum of lengths of all virtual links formed). With the respect to such optimal ranking schemes, it is not clear how close the SN and FN schemes are to being optimal; they are, however, simple and intuitive policy choices that require no optimization in terms of pairing of links for entanglement swapping, and hence are practical for near-term implementation. We defer the discussion of other policy alternatives and comments on the optimality of SN and FN swap-asap to Appendix E.
Let us now compare our policies with the doubling policy, described in Sec. IV.1.
Figure 17: Average waiting time (left) and average age of the youngest end-to-end link (right) for a repeater chain with $n=9$ nodes (eight elementary links), according to the FN, SN, and random swap-asap policies, compared to the SN doubling policy. For small values of $p_{\ell}$, both the FN and SN swap-asap policies outperform the doubling policy in terms of waiting time. The SN swap-asap policy outperforms doubling for the average age for all values of $p_{\ell}$ up to $p_{\ell}=0.5$.
Appendix B Noise model and decoherence
@@
In order to connect with the results presented in the main text, for a given value of $m^{\star}$ as presented in the main text, let us take
@@
Appendix C Entanglement distillation with the BBPSSW protocol
@@
Note that these formulas apply only when the input states to the distillation protocol are so-called isotropic states, given by $f\Phi^{+}+(\frac{1-f}{3})(\Phi^{-}+\Psi^{+}+\Psi^{-})$, with $f\in[0,1]$ being the fidelity. Also, distillation should be considered useful only when it can improve upon the fidelity of the best of the two links. In other words, in order for distillation to be useful, we should have $F_{\text{distill}}(f_{1},f_{2})>\max\{f_{1},f_{2}\}$. In Fig. 18, we plot the $(f_{1},f_{2})$ region defined by this inequality.
From Ref. [79], we have the following relations, telling how the Bell states $\Phi^{\pm}$ and $\Psi^{\pm}$ are transformed by the BBPSSW protocol.
@@
where in the last line we applied the isotropic twirling map $X\mapsto\mathrm{Tr}[\Phi^{+}X]\Phi^{+}+\frac{1-\mathrm{Tr}[\Phi^{+}X]}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-})$ [107, Example 7.25], which is required in order to use the BBPSSW protocol recursively. Therefore, after normalization, the state is
$F_{\text{distill}}(f(m_{1}),f(m_{2}))\Phi^{+}+\frac{1-F_{\text{distill}}(f(m_{1}),f(m_{2}))}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-}).$ (33)
@@
Figure 18: Usefulness of distillation with the BBPSSW protocol [79]. The dark blue shaded region indicates the values of the initial fidelities, $f_{1}$ and $f_{2}$, for which the fidelity after distillation exceeds the fidelity of the best of the two links, i.e., when $F_{\text{distill}}(f_{1},f_{2})\geq\max\{f_{1},f_{2}\}$.

### v3

@@ Introduction
Let us now address the following question. Given a particular number $n_{ch}$ of channels for multiplexing, and given other physical parameters corresponding to channel losses, quantum memory coherence times, etc., what is the optimal number of repeater nodes for a linear quantum repeater chain spanning a certain distance, and what would be the rate and fidelity of end-to-end entanglement distribution for this optimal setting? Intuitively, we do expect such an optimal number of nodes to exist, because as the number of nodes increases, even though the individual links become smaller and thus $p_{\ell}$ and $m^{\star}$ both increase, the non-deterministic nature of entanglement swapping and/or the “age addition” rule, will at some point begin to adversely affect the waiting time as the number of nodes increases. Furthermore, in a (source-)rate limited setting, $m^{\star}$ also saturates to a maximum value with decreasing elementary link length.
@@
Figure 10: Performance of the proposed experimental implementation using our multiplexing policies. Average rate of end-to-end entanglement distribution (a,b), including classical communication overheads, and average fidelity of the end-to-end link (c,d) for entanglement distribution along a 100 km repeater chain with five channels with rare-Earth ($\text{Pr}^{3+}$ ion) (a,c) and diamond vacancy (b,d) based memories using the FN swap-asap policy. The dotted (solid) lines represent $p_{sw}=0.5(p_{sw}=1)$. Both state-of-the-art and near-term values for the fidelity $f_{0}$ of fresh entangled states generated (absorbed) by memory platforms are shown. Using three to five repeaters between the end nodes seems to be optimal for both memory platforms in terms of waiting times.
@@
To answer the question posed above, let us set the end-to-end distance to be 100 km. We also use the FN swap-asap policy for illustrative purposes, and we also include the classical communication overheads, but we leave out distillation-based policies. The physical parameters are chosen to reflect the proposed experimental implementation above. They are as follows:
  • Number of multiplexing channels $n_{ch}$: We choose $n_{ch}=5$, since although the SPDC spectrum in Fig. 9(b) can be divided into nine channels, both the ebit rates of polarization entangled states and their fidelity with respect to the maximally entangled Bell states fall as one moves away from the center of the spectrum.
  • Source rate $R$: We choose $R=5000$ ebits/s. Since the original SPDC source is divided into different frequency channels, the number of ebits per channel is at least lower by a factor of $n_{ch}$. Furthermore, as shown in Fig. 9(b) the sinc function of the SPDC spectrum makes the ebit rates fall as one moves away from the central channel. Thus, to make a conservative estimate, we choose $R$ to be rate corresponding to the dimmest channel. This also allows us to choose the same rate for all channels (consistent with our simulation assumptions that all channels are identical).
  • Fidelity of state at source $f_{s}$: We choose $f_{s}=0.93$, as obtained in the experiment mentioned above. Recall that the fresh elementary link fidelity is given by $f_{e}=f_{s}f_{0}^{2}$.
@@
Since quantum memories and entanglement swapping have not yet been implemented in the experiment, we choose two paradigmatic values for the swapping success probability, which are $p_{sw}=0.5$ (linear optics) and $p_{sw}=1.0$ (perfect gates on solid-state memories). For the quantum memories, we choose two quantum memory platforms for our analysis, a rare-Earth metal based memory ($\rm{Pr}^{3+}$ ions) and a diamond vacancy based memory. The state-of-the-art physical parameters for both the Rare-Earth and Diamond-based quantum memories are taken from Ref. [103, Table 1]. Given the above SPDC source and memory parameters, we can calculate all the relevant simulation parameters following the prescription given above. We plot the average rate and fidelity of end-to-end entanglement distribution both for near-term and state-of-the-art parameter values in Fig. 10. We tabulate the values for the fidelity and entanglement distribution rate for the two memory platforms in the optimal setting, along with the hardware parameters, in Table 1. The optimal number of nodes for the near-term (state-of-the-art) settings in terms of the waiting time seems to be around five to seven (two to four) for both memory platforms, which have $T_{2}$ times on the order of a few milliseconds. In the rate optimal setting (number of nodes), we anticipate entanglement distribution rates for the near-term (state-of-the-art) parameters ranging from a few Hertz to tens of Hertz, with an average end-to-end link fidelity of 60-70% (53-55%). We also see that as the swapping success probability is increased, especially close to the optimal choice of number of repeaters, there is a substantial increase in the rates, although the end-to-end fidelity does not substantially improve. This is also an artifact of choosing the farthest neighbor policy, which prioritizes forming long links. Also, the results of both platforms indicate that the end-to-end link fidelity decreases with an increasing number of nodes.
@@
Since the elementary links have (nearly) unit fidelity for the rare-Earth metal-based memory, distillation is detrimental as shown by the results in Fig. 8. For the diamond vacancy-based memory with state-of-the-art initial state fidelity of 89%, we also find that for the network parameters under consideration, distillation is in fact detrimental. This observation can be attributed to the higher CC overheads compared to the small increase in fidelity that distillation offers in this case. For example, the maximum rates available (for optimal number of repeaters) with deterministic entanglement swapping and using FN-based distill-swap policy is around 1 Hz, which is at least five times smaller than the rates available without distillation.
State-of-the-art parameters
Memory $T_{2}$ $\eta$ $f_{e}$ $p_{\ell}$ $m^{\star}$ $m_{0}$ $R$ $f$
Rare-Earth 1.2 ms 0.69 0.89 0.023 7 1 53 Hz 0.55
Diamond 13 ms 0.5 0.74 0.003 52 22 10 Hz 0.54
Near-term parameters
Rare-Earth 1.2 ms 0.69 0.91 0.046 7 0 190 Hz 0.59
Diamond 13 ms 0.5 0.91 0.04 78 9 342 Hz 0.59
Table 1: Optimized figures of merit for a repeater chain using frequency multiplexing and quantum memories. Hardware parameters for two quantum memory platforms (rare-earth ion and diamond vacancy) [103, Table 1][104, 105], and their respective optimal entanglement distribution rates ($R$) and end-to-end fidelities ($f$), for a 100 km repeater chain, based on the data in Fig. 10. $T_{2}$ is the coherence time of the memory, $\eta$ is the absorption efficiency, i.e., Debye–Waller factor (DWF). in the case of Rare-Earth memory, for emissive memories like the Diamond vacancy-based memory an effective DWF of 0.5 is assumed. $f_{e}$ is the fidelity of a fresh elementary link (which includes the fidelity of the SPDC source state $f_{s}$ and the initial entanglement generated or absorbed by the quantum memories $f_{0}$). For the near-term case we consider $f_{0}=0.99$. Since, solid-state memories are considered, we assume swapping to be deterministic. Also, shown are the parameters $p_{\ell}$, $m^{\star}$, and $m_{0}$ of our theoretical model.
Conclusions
@@
In this paper, we have taken steps in these directions. We have presented practical, quasi-local multiplexing-based policies for long-distance entanglement distribution using quantum repeaters with multiple memories. We call these policies farthest neighbor (FN) swap-asap and strongest neighbor (SN) swap-asap, adapting swap-as-soon-as-possible (swap-asap) policies to multiplexing-based linear networks. These policies go beyond fully local policies, such that the nodes have knowledge of states of the other nodes in the chain, but not necessarily full, global knowledge.
@@
Throughout this work, we have also considered average values of the waiting time and fidelity of the end-to-end link. The behavior of the higher moments of these quantities, and in particular the distributions of these quantities, is an important consideration that has an impact on how well the average estimates the real behavior. We anticipate that this will involve new techniques, and is an interesting direction for future work.
Methods
Noise model and decoherence
@@ Methods
In order to connect with the results presented in “Results”, for a given value of $m^{\star}$ as presented there, let us take
@@
Entanglement distillation with the BBPSSW protocol
@@
Note that these formulas apply only when the input states to the distillation protocol are so-called isotropic states, given by $f\Phi^{+}+(\frac{1-f}{3})(\Phi^{-}+\Psi^{+}+\Psi^{-})$, with $f\in[0,1]$ being the fidelity. Also, distillation should be considered useful only when it can improve upon the fidelity of the best of the two links. In other words, in order for distillation to be useful, we should have $F_{\text{distill}}(f_{1},f_{2})>\max\{f_{1},f_{2}\}$. In Fig. 11, we plot the $(f_{1},f_{2})$ region defined by this inequality. General restrictions on useful entanglement distillation protocols can be found in Ref. [110].
From Ref. [83], we have the following relations, indicating how the Bell states $\Phi^{\pm}$ and $\Psi^{\pm}$ are transformed by the BBPSSW protocol.
@@
where in the last line we applied the isotropic twirling map $X\mapsto\mathrm{Tr}[\Phi^{+}X]\Phi^{+}+\frac{1-\mathrm{Tr}[\Phi^{+}X]}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-})$ [111, Example 7.25], which is required in order to use the BBPSSW protocol recursively. Therefore, after normalization, the state is
$F_{\text{distill}}(f(m_{1}),f(m_{2}))\Phi^{+}\\ +\frac{1-F_{\text{distill}}(f(m_{1}),f(m_{2}))}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-}).$ (33)
@@
Figure 11: Usefulness of distillation with the BBPSSW protocol. The dark blue shaded region indicates the values of the initial fidelities, $f_{1}$ and $f_{2}$, for which the fidelity after distillation with the BBPSSW protocol [83] exceeds the fidelity of the best of the two links, i.e., when $F_{\text{distill}}(f_{1},f_{2})\geq\max\{f_{1},f_{2}\}$.
Classical communication costs for local, quasi-local and global policies
In this section, we aim to discuss some of the details related to our model for calculating the classical communication costs for different policies. As was mentioned in “Results”, based on the amount of information available to nodes when making policy decisions, policies can be classified as local, quasi-local, and global. We remind the reader that in our model, for all policies, no CC is allowed within one MDP step. Consequently, if at the beginning of an MDP step it is determined that multiple interdependent swaps or distillation attempts are to be made, the failure of one leads to the failure of all dependent swaps or distillation attempts.
Figure 12: Classical communication costs of entanglement swapping. (a) A generic network evolution trajectory. The CC cost is estimated in our model as the length of the longest link involved in a swap in a given time-step. The swap at node 6 leads to a CC cost of 3 units, where 1 unit is equal to the heralding time for the elementary link. (b) Evolution of a section of the chain with no connections initially. The end nodes stop waiting for CC if there is a gap in the stream of CC packets they receive.
Consider the example of a generic network state as shown in Fig. 12(a). We ignore distillation in this discussion, the CC cost for which is much easier to estimate and has been explained in “Results”. Let us focus on the actions of node 6 and isolate one of the channels between the nodes of the chain. Node 6 knows at the beginning of the MDP step that it is connected to node 3. Now it will try to generate an elementary link with node 7. As soon as such a link is generated, it will perform a swap. Some of the time evolution trajectories that can emerge are as follows.
To elaborate on such a scheme further, we give another example in Fig. 12(b). We consider a portion of the network in which there are no links to begin with and show how a connected segment of the chain is established, with each node aware of its connections.
At the beginning of the protocol, each node tries to generate an elementary link. Suppose all of these attempts succeed. Now, each node will attempt swaps simultaneously, because the information about their respective elementary link generation success is available in one time step (heralding time). Now, if one of the swaps fails all of them must fail, and this information will be available at each node, allowing them to restart their memories in the next time step. This will also mark the end of one round of the protocol for this portion of the chain and no extra CC will be required. The other possibility is that all the swaps have succeeded, in which case the internal nodes will be free to restart without the need for any CC. They will just send their measurement outcomes to their next node, that is the node they were connected to at the beginning of the outcome, this CC signal will now be relayed along the chain with more information being added at each node until all this information reaches both the end nodes (see Fig. 12(b)). One of the end nodes can keep utilizing this classical information to perform the local operations necessary to create an end-to-end entangled link. (Which one? We can have a set convention that the node on the left, i.e., one receiving CC signals from the right of it, will do Pauli rotations.) As more and more CC has reached them, they get informed about the success and failure of swaps that occurred farther and farther away, and they get to know that they now share a link with some far-off node that has a certain age. The end nodes will stop waiting for more CC once there is a gap of one time step in the CC packets reaching them. This will indicate that they have received all CC from the connected region of the chain.
Error estimates for Monte-Carlo simulations
All simulation results in this work were produced by performing Monte Carlo simulations. Therefore, all average waiting time and fidelity (end-to-end link age) values have some statistical errors. For most hardware parameter regimes considered in the work, averages of the figures of merit were obtained by performing 20 batches of 1000 runs each. One run is considered as starting from a fully disconnected chain and evolving the network until at least one end-to-end link is generated. The average value reported is the mean of means of all the 20 batches. The standard deviation of this mean (of means) falls with increasing number of batches. The choice of number of batches was made to keep the standard deviation less than 5% of the reported mean values. Therefore, statistical error (due to finite statistics generated by the Monte Carlo simulations) is less than 5% for all values reported in our work. Here we would also like to point out that higher moments of the waiting times and fidelities are also an important figure of merit, e.g., if the variance of the waiting time is very large, having a small average does not guarantee that most runs will take a small time. The full probability distributions can also be obtained using our simulations, but this was not the focus of our work. Existing works have looked at these distributions in detail, for example Ref. [26].
Data availability
The data generated through the simulations performed in this work and the experimental data that support the findings of this study are available on request from the corresponding authors.
Code availability
The custom code developed to perform the simulations in this work that support the findings of this study are available on request from the corresponding authors.
Author contributions
SH, SK, HL, and PB conceived the main idea of the work. SK, PB, and SH developed the theory and performed the computations. SK and SH performed the analytical calculations. XC and KC performed the experiments supervised by CW. BK encouraged SH to investigate the role of classical communication. SK, BK, CW, and HL supervised the findings of this work. All authors discussed the results and contributed to the writing of the manuscript.
Competing interests
The authors declare no competing interests.
Acknowledgements.
Supplementary Note 1: Properties of our multiplexing policies
Figure 13: Performance of our multiplexing policies against parallel swap-asap. Average waiting time and average age of the youngest end-to-end link for a repeater chain with $n=7$ nodes, according to the multiplexed (SN and FN swap-asap) and non-multiplexed (parallel swap-asap) policies. The multiplexed versions can enormously reduce average waiting time compared to parallel swap-asap, and at the same time increase end-to-end link fidelities. SN is the better choice for fidelity enhancement, while FN reduces the waiting time more significantly.
In Fig. 13, and for all of the plots shown in this paper, the parameter values that we choose are based on what is computationally feasible, but they are also based on what we expect with actual hardware, as we outline in the main text. The choice of network parameters such as $p_{\ell},m^{\star}$ etc. in specific examples is also motivated by the intent to show the most prominent and overarching trends.
Figure 14: Comparison among different multiplexing policies. Average waiting time and average age of the youngest end-to-end link for a repeater chain with $n=7$ nodes, according to the FN, SN, and random swap-asap multiplexing policies. SN and FN both outperform the random policy, indicating the benefit of using knowledge of the network state. SN is the better choice for fidelity enhancement, while FN reduces the waiting time more significantly.
Figure 15: Multiplexing can beat the detrimental effect of lower link success probability. Heat map of average waiting time (left) (denoted $wt\right>$) and ratio of the average age of youngest end-to-end link to the cutoff (right) as a function of elementary link success probability and number of channels. Contours of fixed values run along diagonals from the top-left to the bottom-right, showing that lower link probabilities can be compensated for by increasing the number of channels. Both the average waiting time and the average age of the youngest end-to-end link reduce as the number of channels increases, showing the usefulness of multiplexing. For the waiting time results we use the FN policy, and for the average age results we use the SN policy.
Figure 16: Multiplexing can beat the detrimental effect of lower memory coherence time.Heat map of average waiting time (left) (denoted by $wt\right>$) and ratio of the average age of youngest end-to-end link to the cutoff (right) as a function of link cutoff age and number of channels. Contours of fixed values run along diagonals from the top-left to the bottom-right, showing that lower cutoffs, which translates to memories with smaller coherence times, can be compensated for by increasing the number of channels. For waiting time results we use the FN policy and for average age results we use SN policy.
Of course, FN, SN, and random are not the only three possible options for multiplexing-based policies. For example, one could develop a ranking scheme based on a combination of age and distance, or one could perform a search over all possible pairings of links, looking for the optimal pairings in terms of some figure of merit (such as the sum of lengths of all virtual links formed). With respect to such optimal ranking schemes, it is not clear how close the SN and FN schemes are to being optimal; they are, however, simple and intuitive policy choices that require no optimization in terms of pairing of links for entanglement swapping, and hence are practical for near-term implementation. We defer the discussion of other policy alternatives and comments on the optimality of SN and FN swap-asap to Supplementary Note 4.
Let us now compare our policies with the doubling policy, described in the main text.
Figure 17: Performance of multiplexing policies compared to the doubling policy. Average waiting time (left) and average age of the youngest end-to-end link (right) for a repeater chain with $n=9$ nodes (eight elementary links), according to the FN, SN, and random swap-asap policies, compared to the SN doubling policy. For small values of $p_{\ell}$, both the FN and SN swap-asap policies outperform the doubling policy in terms of waiting time. The SN swap-asap policy outperforms doubling for the average age for all values of $p_{\ell}$ up to $p_{\ell}=0.5$.
Supplementary Note 2: Entanglement distillation with the BBPSSW protocol for more than two links
Recall that if $f_{1}$ and $f_{2}$ are the fidelities of the two links being distilled using the BBPSSW protocol, then the success probability $P_{\text{distill}}(f_{1},f_{2})$ of the protocol and the resulting fidelity $F_{\text{distill}}(f_{1},f_{2})$ are given by